# MCTR: Multi Camera Tracking Transformer

### v1

@@ 4.2 AI City Challenge dataset
Table 2: Results on AI City Challenge dataset
HOTA IDF1 MOTA
Env. MCTR MOTR MCTR MOTR MCTR MOTR
S002 62.22 66.08 63.18 58.39 81.73 92.58
S004 70.38 61.16 78.48 54.43 94.47 96.24
S006 82.64 76.27 89.79 69.19 95.53 96.98
S007 51.23 51.99 53.36 41.96 90.15 93.04
S010 68.14 64.27 75.39 54.36 94.96 94.48
HOTA IDF1 MOTA
Env. MCTR MOTR MCTR MOTR MCTR MOTR
S011 46.63 40.40 50.19 26.57 96.22 96.48
S012 60.37 55.37 60.63 42.95 97.81 95.13
S015 56.06 45.17 57.47 32.60 90.65 93.62
S016 54.03 53.55 56.46 40.92 92.79 96.55
S019 46.49 38.70 47.31 25.36 96.63 96.04
@@
Table 2 shows the results for MCTR and MOTR on the AI City Challenge dataset. MCTR again outperforms MOTR on IDF1 and HOTA metrics indicating improved long-term tracking and robustness to occlusions. MOTR on the other hand tends to have better MOTA score, mainly due to a higher detection precision and recall. The improved detection performance of MOTR is likely due to their use of deformable DETR [56] which is known to perform better than standard DETR at detecting small objects, and to a prevalence of highly occluded persons (e.g. only head is visible behind shelves) in the dataset leading to a significant number of small objects.
@@ 5 Conclusion & Limitations
Compared with highly optimized heuristic systems that achieved state of the art performance in the MMP-Tracking competition, the MCTR does lag behind, especially in terms of IDF1 score. One reason for the lower long-term tracking performance might be the Markovian nature of the temporal component in MCTR. Track embeddings are updated using only information from the current frame, and the track embeddings from the previous frame. Early attempts to use a longer temporal history did not result in improved performance. Finding better ways to incorporate more temporal information and using motion information may lead to improved long term tracking performance.
Another limitation of MCTR, and of end-to-end systems in general, is the requirement for large amounts of labeled data necessary for training. In the case of MCTR, this limitation is even more stringent, as the learned model is dependent on the particular camera setup it is trained on due to the camera-view specific cross-attention modules in the tracking module. An interesting question is whether a more general, camera setup independent, end-to-end model can be devised. This could potentially be achieved by using camera calibration parameters as inputs to adapt a common cross-attention module to each camera view, rather than using separate cross-attention modules. We leave this direction for future work.
Appendix A Model Parameters
The DETR model is the same as in [7], with resnet50 as a backbone, six encoding and six decoding layers, twenty 256 dimensional detection queries, eight head attention layers and 2048 dimensional feed-forward layers. The tracking module also uses twenty 256 dimensional tracking queries, eight head attention layers and 2048 dimensional feed-forward layers. Training uses the Adam optimizer [21], with a starting learning rate of 1e-4 which decays by a factor of 0.3 every 20 epochs. The model is trained with short video segments (see Section 3.5) for 30 epochs, followed by training with increasingly longer segments for 100 epochs. Training was done on a machine with single A100 GPU with 500GB CPU memory and 64 CPU cores. It takes between 16-24 hours for entire training run per scene depending on the size of the training dataset for individual scenes.

### v2

@@ 4.2 AI City Challenge dataset
Figure 9 shows the results for MCTR and MOTR on the AI City Challenge dataset.  [^4] We do not run ReST on this dataset because of its poor performance on MMPTrack, and the lack of ground truth camera calibration parameters which are required by ReST. MCTR again outperforms MOTR on IDF1 and HOTA metrics indicating improved long-term tracking and robustness to occlusions. MOTR on the other hand tends to have better MOTA score, mainly due to a higher detection precision and recall. The improved detection performance of MOTR is likely due to their use of deformable DETR [58] which is known to perform better than standard DETR at detecting small objects, and to a prevalence of highly occluded persons (e.g. only head is visible behind shelves) in the dataset leading to a significant number of small objects.
@@ 5 Conclusion & Limitations
Compared to highly optimized heuristic systems that led the MMP-Tracking competition, MCTR does lag behind, particularly in IDF1 score. The lower long-term performance may stem from MCTR’s Markovian approach, where track embeddings are updated only with current and previous frame information. Early attempts to use longer temporal history did not improve results. Finding better ways to integrate more temporal and motion information could enhance long-term tracking performance.
Another limitation of MCTR, and of end-to-end systems in general, is the need for extensive labeled data for training. Moreover, MCTR’s reliance on camera-view specific cross-attention modules means it is closely tied to the particular camera setup used during training. An interesting question is whether a more general, camera setup independent, end-to-end model can be devised. This could involve using camera calibration parameters to adapt a single cross-attention module for all views, rather than separate modules. We leave this direction for future work.